OpenStudy (anonymous):

How would you simplify 3 times the square root of 72

OpenStudy (anonymous):

Find The Square Root.

OpenStudy (anonymous):

3 sqrt 72 = 3 sqrt (36 • 2) = 3 sqrt 36 sqrt 2 = 3(6) sqrt 2 = 18 sqrt 2

OpenStudy (cwrw238):

look for factors of 72 whcih are exact squares 72 = 2 * 36 3 * sqrt36 * sqrt2 = 3*6*sqrt2 = 18 sqrt2

OpenStudy (jhannybean):

\[\large 3\sqrt{72}\]\[\large 3 \sqrt{ \color{green}{6 \cdot 6} \cdot2}\]\[\large 3(\color{green}{6})\sqrt{2}\]\[\large 18\sqrt{2}\]